OpenStudy (baby456):

look at the three sequense below. sequence 1 F(0)=45 f(n)=F(n-1)-4 sequence 2 f(n)=45-4n sequence 3 45, 41, 37, 33 ,29, which of the above sequences are arithmetic ? a.sequence 1 b.sequence 3 c. sequence 1 and 3. d. sequence 1, 2, and 3 explain and i will fan and medal!

OpenStudy (amoodarya):

f(n)=F(n-1)-4 is arithmetic because f2-f1=-4 and f3-f2=-4 and so on f(n)=45-4n is arithmetic too because f(n)-f(n-1)=45-4n-(45-4(n-1))=-4n+4n-4=-4

OpenStudy (amoodarya):

45, 41, 37, 33 ,29 45-41=4 41-37=4 37-33=4 33-29=4 it is arithmetic too